OpenStudy (anonymous):

Kyle is 52 inches tall. The distance from the top of his head to the tip of his shadow is 60 inches. Approximately how long is Kyle's shadow? A. 30 inches B. 40 inches C. 79 inches D. 102 inches

OpenStudy (wolf1728):

You are given 2 sides of a right triangle: Distance from top of head to shadow = 60" = hypotenuse Height = 52" = 1 side So 60² = 52² + side² side² = 60² - 52² side² = 3,600 -896 side² = 2,704 side = 29.93326 (which is the length of the shadow)

OpenStudy (anonymous):

Thanks! That's the right answer!